Find the slope-intercept form for the line perpendicular to 2x –3y=-6 and passing through (4,–9)


1
Expert's answer
2021-11-08T17:34:00-0500

For perpendicular lines, "m_1m_2 = -1"


From "2x-3y=-6" , we have that "y=\frac{2}{3}x+2"


Let "m_1=\frac{2}{3}"

=> "m_2=\frac{-3}{2}"


"m_2=\frac{y-y_1}{x-x_1}"



"\frac{-3}{2}=\frac{y-(-9)}{x-4}=\frac{y+9}{x-4}"


=> "-3(x-4)=2(y+9)"

"12-3x=2y+18"


"=> y = \frac{-3}{2}x-3"